A rocket is fired at a speed of 70.3 m/s from ground level, at an angle of 61.7° above the horizontal. The rocket is fired toward an 11.4 m high wall, which is located 28.2 m away. By how much does the rocket clear the top of the wall?
Answer #1

Velocity = 70.3 m/s at 61.7 °

Horizontal velocity vh = 70.3 cos 61.7 =33.33 m/s

vertical velocity vv = 70.3 sin 61.7 = 61.9 m/s

Horizontal distance of wall = 28.2m

time to reach wall = 28.2/33.33 = 0.84 s

vertical distance at 0.84 s = 61.9 * 0.84 - 0.5 * 9.8 * 0.842 = 48.54 m

So rocket clears wall by 48.54 -11.4 = 37.14 m